### v2

# MSAGSM: Multi-Scale Attention Gated Shifting for Precise Event Spotting

Abstract
Precise Event Spotting (PES) in sports videos requires frame-level recognition of fine-grained actions from single-camera footage. Existing PES models typically incorporate lightweight temporal modules such as the Gate Shift Module (GSM) or the Gate Shift Fuse to enrich 2D CNN feature extractors with temporal context. However, these modules are limited in both temporal receptive field and spatial adaptability. We propose a Multi-Scale Attention Gate Shift Module (MSAGSM) that enhances GSM with multi-scale temporal shifts and channel grouped spatial attention, enabling efficient modeling of both short and long-term dependencies while focusing on salient regions. MSAGSM is a lightweight, plug-and-play module that integrates seamlessly with diverse 2D backbones. To further advance the field, we introduce the Table Tennis Australia dataset, the first PES benchmark for table tennis containing over 4,800 precisely annotated events. Extensive experiments across four PES benchmarks demonstrate that MSAGSM consistently improves performance with minimal overhead, setting new state-of-the-art results. All code and datasets are publicly available at https://anonymous.4open.science/r/MSAGSM-B0A4.
1 Introduction
Detecting the precise moment when events occur in a video is a core challenge in video understanding, particularly in the single camera setting commonly used in most sports recordings. Precise Event Spotting (PES) aims to localize and classify events at frame-level accuracy, typically within a 1–2 frame tolerance [12]. This fine-grained precision is critical in sports scenarios, where even a few frames of deviation can correspond to a different event entirely (see Figure 1). Accurate PES provides a strong foundation for downstream tasks such as ball tracking, player positioning, and automatic highlight generation. Moreover, precise spotting plays a vital role in post-game analysis by highlighting key moments and reducing the need to manually browse full-length videos. For instance, according to Table Tennis Australia’s sports analytics team, annotating events in a single match video can take 4–5 hours. Automating this process not only streamlines analysis but also enables timely feedback for athletes and coaches.
Recent advancements in video understanding have explored various approaches to this task, including transformer-based models [1, 54], multimodal methods [38, 43], and self-supervised frameworks [5]. Among these, E2E-Spot [12] has emerged as a strong baseline, inspiring several subsequent models such as the current state-of-the-art (SOTA) T-DEED [44], UGLF [37], and ASTRM [27]. A key component in E2E-Spot is the Gate Shift Module (GSM) [31], which enables 2D CNNs to learn temporal dynamics by shifting features across adjacent frames. While effective, GSM is limited in its temporal range and does not explicitly model longer-term dependencies.
Despite these advancements, PES remains a highly challenging task due to its unique characteristics. Events are often difficult to distinguish from adjacent frames, sometimes even for human annotators without access to extended temporal context [44]. Moreover, different events may require varying amounts of temporal information to be correctly identified, further complicating model design. In addition to temporal reasoning, effective PES also demands accurate spatial attention: in sports videos, a large portion of each frame typically consists of background, making it essential for models to focus on small but informative regions such as the ball or a player’s hand position to make correct predictions. Another practical limitation lies in the availability of suitable datasets. Most publicly available sports video datasets are derived from broadcast footage of elite sports [12, 11, 9, 48], which are often professionally edited and filmed from ideal angles. However, in many real-world sports settings, such as local or developmental leagues, matches are recorded using handheld devices with varied perspectives, inconsistent quality, and frequent occlusions.
Figure 1: Example of Precise Event Spotting in table tennis: the exact moments when the player contacts the ball during a serve (highlighted in red) and when the ball bounces on the table (highlighted in yellow). The ball locations are circled accordingly. t represents the timestamp at the frame level.
To address these challenges, we propose the Multi-Scale Attention Gate Shift Module (MSAGSM), a lightweight extension of GSM that overcomes its limitation of accessing only adjacent frames by incorporating multi-scale temporal shifts. To mitigate the noise introduced by larger temporal windows, we further design a channel-grouped spatial attention mechanism that learns spatial masks within channel groups to guide shifting, differing from transformer multi-head attention by avoiding costly query–key–value interactions. This design enables MSAGSM to capture longer-range dependencies while focusing on salient regions (e.g., ball or player positions), and it can be seamlessly integrated into 2D CNN backbones with minimal overhead compared to 3D models.
In addition, we introduce a new dataset for PES: the Table Tennis Australia (TTA) dataset, recorded and curated in collaboration with Paralympics Australia’s sports analysts, providing more realistic recordings compared to existing datasets. The dataset includes 39 games and a total of 4,878 annotated events across eight event types: serve, bounce, forehand, and backhand—each distinguished between the near and far sides of the table. To our knowledge, TTA is the first PES dataset specifically designed for table tennis. Compared to existing PES datasets, TTA is both sizable and challenging. Its scale is comparable to FineGym [28] and surpasses that of FigureSkating [11]. Moreover, unlike tennis [12], table tennis involves more densely packed, fine-grained events within shorter time spans, presenting a more demanding benchmark for PES.
Our main contributions are summarized as follows: (1) proposing MSAGSM, a lightweight module that combines multi-scale temporal shifts with spatial attention, and can be efficiently plugged into existing 2D CNN backbones; (2) introduce the TTA dataset, the first PES dataset for table tennis, containing 39 games and 4,878 annotated events across eight classes; and (3) show that MSAGSM, when integrated with standard 2D backbones, achieves SOTA performance across multiple PES benchmarks, including TTA.
Figure 2: Overview of our proposed architecture for Precise Event Spotting. The input video is processed through a 2D CNN backbone for spatial feature extraction, followed by the Multi-Scale Attention Gate Shift Module (MSAGSM) for spatio-temporal modeling. MSAGSM combines channel-group spatial attention and multi-scale temporal shifting to capture varying temporal dependencies. The resulting features are passed through a GRU-based temporal model to predict event class probabilities at the frame level.
2 Related Work
Spatial-Temporal Modelling: Significant progress has been made in video understanding tasks, such as video classification and action recognition, driven by the development of deep CNNs and Transformer-based architectures [6, 7, 20, 2, 35]. One major challenge in these models is their high memory and computational requirements, which make them less suitable for deployment on resource-constrained devices compared to lightweight 2D CNNs.
To address this, recent research has focused on developing efficient and effective temporal modeling techniques built on top of 2D CNNs, minimizing computational complexity while retaining performance. Several approaches attempt to decompose 3D convolutions into separate spatial (2D) and temporal (1D) components [24, 36, 45, 14]. Others propose hybrid models that combine 3D CNNs with 2D CNNs to reduce overall cost [22].
A key advancement in this direction is the Temporal Shift Module (TSM) [17], which enables 2D CNNs to capture temporal information by shifting a portion of the feature channels along the temporal dimension. This parameter-free operation significantly improves efficiency over 3D or 2D+1D methods and demonstrated state-of-the-art performance in action recognition at the time. However, TSM’s shift mechanism is manually hardcoded and does not adapt to the input content.
To address this limitation, the GSM [31] was introduced. Built upon TSM, GSM employs an additional spatio-temporal 3D convolutional kernel as a gating mechanism to selectively decide which features to shift forward or backward based on contextual information. Its variant, Gate Shift Fuse (GSF) [32], adds a point-wise convolution block before and after the gating operation to enable interaction between spatial and spatio-temporal feature groups, further enhancing modeling capability.
Precise Event Spotting: In video understanding tasks such as video classification and action recognition, a common approach is to sparsely sample frames from the entire video [8, 39, 15]. In contrast, PES requires dense frame sampling, particularly for sports videos due to the fast motion and high frequency of events occurring within short time spans. Compared to Temporal Action Detection, which focuses on predicting temporal intervals of actions [18, 30], PES aims to identify actions using a single key frame. This is especially relevant in sports scenarios, where the start and end boundaries of actions (e.g., a forehand shot) can be visually ambiguous and difficult to define precisely [46].
To address this ambiguity, the SoccerNet dataset introduced the task of Action Spotting [9], which focuses on identifying key frames of important events within a relaxed temporal window (e.g., $\pm$50 frames). Subsequently, PES was proposed to enforce a stricter temporal tolerance, requiring models to be accurate within just a few frames.
E2E-Spot [12] serves as a strong PES baseline, combining RegNetY [25] with the Gate Shift Module (GSM) [31] for temporal modeling. Building on this, T-DEED [44] addressed the challenge of distinguishing between adjacent frames with highly similar visual appearance by incorporating the Scalable Granularity Perception module [29] to enhance discriminability among temporally close features. UGLF [37] takes a different direction by integrating the vision-language model GLIP [16], marking the first use of vision-language features for PES. Most recently, ASTRM [27] further advanced the field by improving long-range temporal dependency through spatial–temporal feature injection into the feature extractor and proposing the Soft Instance Contrastive (SoftIC) loss to mitigate class imbalance.
Beyond visual features, audio signals can also serve as important cues in sports, for instance, crowd noise after a goal. ASTRA [43], a transformer-based encoder-decoder model designed for soccer, integrates both visual and audio embeddings, using learnable queries in the decoder to jointly attend to multi-modal cues.
Other Sports Video Tasks: Ball tracking methods such as the TrackNet family [13, 33, 4, 26, 47, 19] focus on recovering ball trajectories, with TrackNetV4 introducing motion-aware attention maps from frame differences. While related, tracking localizes objects whereas PES targets the timing and semantics of fine-grained events. These approaches are thus complementary, and accurate trajectories could provide useful cues for future event detection.
Sports Video Datasets: Existing PES datasets are often acquired from broadcast videos from high-level matches such as [9, 28, 48, 12, 11]. These datasets are fine-grained, where all event and class labels relate to a single activity. Other sports-related video datasets include ball tracking [13, 33, 34, 47] and sports video captioning [42, 23, 49, 51].
Despite progress, PES methods still struggle to capture long-range temporal dependencies while attending to fine-grained spatial cues. In addition, most datasets rely on broadcast footage with fixed cameras and ideal conditions, which overlook the challenges of community-level sports. To bridge these gaps, we propose MSAGSM, a lightweight block for enhanced temporal reasoning and spatial focus, and introduce TTA, the first table tennis PES benchmark with consumer-grade courtside videos featuring variable viewpoints, camera motion, and frequent occlusions.
3 Multi-Scale Attention GSM
3.1 Model Overview
Figure 2 illustrates an overview of our proposed MSAGSM architecture. It enables 2D CNNs to capture varying spatio-temporal dependencies through an efficient combination of multi-scale gated shifting and channel-group spatial attention mechanisms. The proposed method consists of two main components: (1) a multi-head channel-grouped spatial attention module, which splits the input features into multiple groups and applies attention independently to each group, allowing the model to focus on spatially informative regions; and (2) a multi-scale GSM, which introduces multiscale temporal dilations to shift features across varying temporal distances. This design allows the model to directly access a broader temporal context compared to the original GSM.
3.2 Channel Grouped Spatial Attention
Our intuition is rooted in the observation that, in sports videos, most critical events are concentrated around key entities such as players, referees, and the ball. In contrast, a large portion of each frame consists of background content that contributes little to event detection. Motivated by the effectiveness of attention mechanisms in prior work [41, 6, 3, 40], we extend the original GSM into a channel-grouped spatial attention GSM. This design allows the model to focus on spatially informative regions while suppressing irrelevant background, thereby improving the accuracy of event spotting. Figure 3 shows the details of the channel-grouped spatial attention mechanism.
We generate spatial attention maps using 2D convolutional layers to highlight the most informative regions within a feature map. Given an intermediate feature tensor, we first split it into multiple groups along the channel dimension, where the number of groups corresponds to a predefined number of attention heads—a tunable hyperparameter.
The motivation behind using multiple spatial attention heads is to account for the varying number of salient regions required across different sports. For example, in tennis, the model must simultaneously attend to multiple entities such as two players and a fast-moving ball, whereas in diving, the focus is typically on a single athlete. Conventional spatial attention mechanisms produce a single attention map, which limits the model’s capacity to flexibly adapt to such variation. In contrast, our multi-head design enables the model to learn multiple independent attention maps, allowing it to focus on different regions as needed, and offering greater flexibility and expressiveness across diverse sporting contexts.
For each head, a 2D convolutional layer generates a spatial attention map based on the original features. These maps are applied independently to their corresponding feature chunks, allowing each head to specialize in distinct spatial regions in a data-driven manner. Finally, the attended feature chunks are concatenated to reconstruct a unified refined feature representation.
Figure 3: Illustration of the spatial attention mechanism in the proposed MSAGSM module. Feature maps from the 2D CNN are divided into multiple channel-wise chunks (e.g., $C/4$ per group). Each chunk is processed independently through a spatial attention module to highlight relevant regions (e.g., players, ball) while suppressing background noise. The attended features are then concatenated and forwarded to the temporal modeling block.
3.3 Multi-Scale GSM
We adopt the GSM as a reference architecture, as it has been proven effective in the PES domain [12, 44, 37]. GSM performs a gating-based temporal shifting mechanism inspired by TSM [17] and GST [22]. However, a key limitation of GSM is that it only performs shifts across adjacent frames, restricting its ability to capture longer range temporal dependencies directly, which is often important in sports videos to classify events and discriminate similar frames.
We propose the MSAGSM, which extends GSM by replacing fixed, adjacent frame shifting with a learnable spatially aware gating mechanism that enables direct access to longer temporal ranges. The architectural design of MSAGSM is illustrated in Figure 4.
Figure 4: Structure of the proposed Multi-Scale Gate Shift Module. Given input features, a 3D CNN generates gate maps corresponding to different temporal shift ranges (e.g., $\Delta t=\pm 1$, $\pm 4$). Features are shifted left and right across multiple scales, gated accordingly, and combined through concatenation and stacking. This design enables the model to adaptively capture both short and long-term temporal dependencies in a data-driven manner, while preserving the original information.
Given a sequence of feature maps $X\in\mathbb{R}^{C\times T\times H\times W}$, where $C$ is the number of channels, $T$ is the temporal length, and $H$ and $W$ denote the height and width of a frame, respectively. For each temporal distance scale, MSAGSM first applies a gating layer implemented as a 3D CNN. This gating layer receives a temporal window of three feature maps sampled across time centered around the current timestamp. Two gating maps, corresponding to the left and right shifts, are generated for each feature map. These gates are passed through the tanh activation function to constrain their values to the range (1,+1)(-1,+1), ensuring stable modulation:
$(G_{\text{left}},G_{\text{right}})=\tanh(\text{Conv3D}(X)).$ (1)
The feature map is split into two groups, $X=[X_{\text{left}},X_{\text{right}}]$, each multiplied element-wise by its corresponding gate map and then temporally shifted left or right via a rolling operation:
$\hat{X}_{\text{left}}$ $=\text{Roll}(X_{\text{left}}\odot G_{\text{left}},-\delta),$ (2)
$\hat{X}_{\text{right}}$ $=\text{Roll}(X_{\text{right}}\odot G_{\text{right}},+\delta).$ (3)
Here, $\odot$ denotes element-wise multiplication, and $\delta$ is the temporal shift distance (in frames).
Padding is applied to maintain shape integrity and prevent feature loss at the sequence boundaries. To preserve important information, the shifted features are added back to the original input via a residual connection. The gated, shifted features are then concatenated along the channel dimension to restore the original shape,
$\hat{X}=[\hat{X}_{\text{left}},\hat{X}_{\text{right}}]+X.$ (4)
Finally, the outputs from different distance levels are stacked along a new dimension and passed through a learnable weighting mechanism. This layer adaptively determines the importance of each temporal scale, enabling the model to incorporate information across varying temporal distances. Formally, let $\{\hat{X}^{(i)}\}_{i=1}^{N}$ denote the outputs from $N$ different shift ranges. The final output is computed as:
$\hat{X}_{\text{final}}=\sum_{i=1}^{N}w_{i}\cdot\hat{X}^{(i)},\quad\text{where }\sum_{i=1}^{N}w_{i}=1,\;w_{i}\geq 0.$ (5)
Here, $w_{i}$ are learnable scalar weights normalized via a softmax function. This design allows the model to emphasize the most relevant temporal scales while introducing only minimal computational overhead compared to the original GSM.
4 Empirical Study
In this section, we describe the datasets (with details on the new TTA dataset), outline implementation settings, present ablation studies, and compare MSAGSM with prior methods and state-of-the-art baselines.
4.1 Datasets
We evaluate our method on four fine-grained sports video datasets with frame-level annotations: Tennis [53], Figure Skating [11], FineDiving [48], and our proposed TTA dataset.
Figure 5: Example frames from the proposed TTA dataset. The top row shows samples from World Para Future matches, while the bottom row displays clips from the 2024 Paris Paralympic Games. The dataset captures diverse viewing angles, frequent occlusions, and high event density, representative of real-world table tennis matches.
The Tennis dataset consists of 3,345 video clips from 28 matches, with frame rates of 25 or 30 FPS. It contains 33,791 frame-accurate annotations spanning six event classes. The Figure Skating dataset includes 11 broadcast videos (all 25 FPS), covering 371 short program performances from major international competitions between 2010 and 2019. It includes 3,674 annotated events across four classes. Following previous work [11, 44, 37], we adopt two evaluation splits: a competition split (FS-Comp) and a performance split (FS-Perf). The FineDiving dataset comprises 3,000 diving clips with temporal segment annotations, totaling 7,010 events. These have been converted into frame-level event labels across four classes using the annotation protocol from [12].
We also introduce the Table Tennis Australia (TTA) dataset details shown in Table 1, the first PES benchmark for table tennis. It consists of 39 full-game videos recorded at 30 FPS, totaling 4,878 annotated events across 8 classes: serve, bounce, forehand, and backhand, each labeled separately for near and far sides of the table. Examples of the TTA dataset are presented in Figure 5. Events are annotated following labeling protocols from Table Tennis Australia. For example, a frame is labeled as an event when the ball contacts the paddle or undergoes significant motion immediately after. In cases of occlusion (e.g., when a player blocks the ball near the camera), temporal context from preceding and following frames is used to estimate the most accurate event frame. For more details, please refer to the Supp. §2.
Table 1: Comparison of PES benchmark datasets. TTA (Bold) offers higher event density and realistic recording conditions.
Dataset Recording Events Classes Density
TTA (Ours) Handheld 4,878 8 7.73
Tennis Broadcast 33,791 6 3.81
Figure Skating Broadcast 3,674 4 1.58
FineDiving Broadcast 7010 4 1.49
Figure 6: Comparison of event density (average number of events) across varying temporal window sizes. TTA shows the highest event density across all ranges, reflecting its fine-grained and fast-paced nature.
Table 2: Performance comparison (mAP %) at different temporal tolerances ($\delta=0,1,2$) across five fine-grained sports video datasets. The best result in each column is shown in bold, and our proposed method is italicized. A $\star$ marks the best result within each model–backbone group. All models are evaluated with a fixed clip length of 100 frames.
Model Temp TTA Tennis FS_Comp FS_Perf FineDiving P (M) / GFLOPs $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 ASTRA[43] - 28.12 54.16 60.95 49.98 88.67 97.49 35.87 73.45 84.71 39.35 78.10 90.05 27.28 65.01 82.74 4.49 / 42.20 ASFormer[50] - 21.68 40.69 61.08 42.57 89.34 93.39 32.60 50.29 55.97 16.34 36.46 58.41 15.86 66.82 87.11 3.21 / 41.94 E2E-Spot[12] GSM 34.36 66.42 76.90 51.42 90.03 96.48 36.44 75.39 89.33 39.32 81.27 94.14 27.31 67.89 86.97 4.45 / 42.18 E2E-Spot GSF 36.17 63.70 72.19 49.68 89.01 96.55 36.25 75.49 89.32 40.07 81.55 94.02 29.41 67.57 86.22 4.45 / 42.20 E2E-Spot Ours 36.43 69.50 78.12 55.51 91.34 96.51 38.34 77.72 90.04 46.54 84.12 96.90 29.83 71.03 87.26 4.51 / 45.06 T-DEED GSM 33.32 60.69 69.30 48.82 89.21 97.15 38.61 71.69 90.02 42.53 79.62 90.36 26.86 66.18 84.02 16.42 / 43.34 T-DEED[44] GSF 26.27 54.04 64.51 50.08 89.58 97.06 38.83 75.52 92.15 38.89 79.82 94.79 28.50 67.04 83.24 16.42 / 43.34 T-DEED Ours 28.06 60.77 72.50 52.76 90.68 96.67 39.06 75.55 87.61 44.66 82.41 94.12 26.88 66.70 84.60 16.48 / 46.20 UGLF[37] GSM 34.36 63.52 71.16 50.31 89.73 97.16 36.99 77.25 90.44 38.29 79.24 93.08 27.11 66.88 84.94 4.45 / 42.18 UGLF GSF 36.17 63.70 72.19 50.94 89.39 97.06 34.84 72.61 86.90 37.41 82.06 94.38 29.81 66.65 84.27 4.45 / 42.20 UGLF Ours 33.93 64.82 74.84 53.10 90.98 97.10 39.66 78.86 91.13 46.59 82.64 96.12 29.59 67.21 85.34 4.51 / 45.06
Unlike existing datasets, TTA videos are collected using handheld cameras by professional analysts from Paralympics Australia. The dataset includes elite level competitions such as the Paralympics and World Para Future. This real-world, single camera setting introduces challenges such as frequent occlusions, motion blur, and dynamic angles—conditions closer to practical applications—unlike the broadcast quality videos used in most prior work. As shown in Figure 6, TTA exhibits the highest event density among all datasets, averaging 7.73 events per 100 frame window, compared to 3.81 in Tennis and significantly fewer in others. This emphasizes the need for models that can capture fine-grained temporal structure and distinguish between visually similar frames within narrow time intervals.
4.2 Evaluation
Since our proposed module, MSAGSM, is designed to be a lightweight and drop in replacement for GSM, we follow the same training procedure as the E2E-Spot baseline [12], replacing GSM with MSAGSM in the backbone. The model performance is evaluated using the mean Average Precision (mAP), computed by averaging the Average Precision across all event classes. We report mAP under $\delta\in\{0,1,2\}$, where $\delta=0$ requires an exact-frame match and $\delta=1/2$ allow a $\pm 1/\pm 2$-frame tolerance.
4.3 Implementation Details
We closely follow the training protocol of E2E-Spot [12], adopting its default configurations for a fair comparison. Specifically, all models are trained with input clips of length $L=100$ frames, a batch size of 8, and 5,000 randomly sampled clips per epoch. Training runs for 50 epochs using the AdamW optimizer [21] with a base learning rate of $1\times 10^{-3}$, 3 warm-up epochs, and a cosine decay schedule. As in E2E-Spot, positive classes in the cross-entropy loss are weighted by $w=5$ to handle class imbalance. We also adopt the same backbones—RegNetY-200 [25] and ResNet-18 [10]—from the original framework. Unless otherwise noted, all experiments are conducted on a single NVIDIA L40S GPU.
4.4 Comparison with State-of-the-Art
To evaluate MSAGSM, we compare against five representative PES models. ASTRA [43] and ASFormer [50] serve as Transformer baselines, while E2E-Spot [12], T-DEED [44], and UGLF [37] are CNN-based methods combining a 2D backbone, temporal module, and decoder. E2E-Spot and UGLF both employ GSM [31], whereas T-DEED adopts GSF [32].
For CNN-based methods, we test both the original temporal module and a drop-in replacement with our MSAGSM, and we also compare different temporal modules within the same backbone–decoder setup. All models are trained from official codebases under identical conditions.
Main Results. Table 2 presents results on five fine-grained sports video datasets under three temporal tolerance thresholds ($\delta=0$, $\delta=1$, $\delta=2$). Across the board, our proposed MSAGSM module consistently outperforms existing temporal modeling approaches. For example, within the E2E-Spot architecture using RegNetY-200 [25], MSAGSM improves upon GSM by +4.09 mAP on Tennis at $\delta=0$, by +3.08 mAP on TTA at $\delta=1$, and by +3.14 mAP on FineDiving at $\delta=1$. Similar performance gains are observed when MSAGSM replaces the original temporal module in more complex models such as T-DEED and UGLF.
Out of all possible comparisons across model backbone configurations, MSAGSM achieves the best result in 37 out of 45 cases when compared directly against other temporal modules. Furthermore, it ranks first in 13 out of 15 columns when considering overall best scores, setting new SOTA performance simply by replacing the temporal component.
Backbone and Module Versatility. To evaluate the generality of our approach, we further experiment with alternative backbones, specifically ResNet-18 [10] in E2E-Spot, T-DEED, and UGLF architectures. For simplicity, the main table includes RegNetY-200 results only, while additional results with ResNet-18 [10] are provided in the Supp. §3.1. For both E2E-Spot and UGLF, which use relatively simple temporal modules, our MSAGSM yields substantial gains with minimal complexity overhead, achieving the best performance in 14/15 and 13/15 evaluation groups, respectively. The improvement is less pronounced for T-DEED, which already employs a more sophisticated temporal design. In this case, adding MSAGSM may introduce redundant or overlapping temporal cues, making feature aggregation less effective.
Efficiency. MSAGSM remains lightweight, requiring only 4.5M parameters and 45.06 GFLOPs in its default setting. This is comparable to existing temporal modules and only slightly larger in size, yet it consistently delivers higher accuracy. Its compact design makes MSAGSM a practical drop-in replacement for both lightweight and large-scale models built on 2D pretrained CNN backbones.
Summary. These results demonstrate the effectiveness of MSAGSM as a robust and efficient temporal modeling module. Its consistent improvements across diverse models and baseline temporal modules, particularly on our proposed TTA benchmark, underscore its suitability for capturing the high-frequency, fine-grained temporal dynamics inherent in sports video understanding.
Table 3: Ablation study on attention head count, temporal distance range, and clip length using the MSAGSM module on the TTA dataset. All results are reported in mAP (%), default clip length=100, best result is in bold.
Setting Value $\delta=1$ $\delta=2$ P(M) / GFLOPs
MSAGSM w/o Attention
E2E-GSM - 66.42 76.90 4.45 / 42.18
Distance [1,2] 66.08 77.31 4.47 / 43.26
Distance [1,2,3] 65.82 74.67 4.49 / 44.32
Distance [1,2,3,4] 64.39 74.91 4.51 / 45.40
GSM with Channel Group Spatial Attention
E2E-GSM - 66.42 76.90 4.45 / 42.18
Heads 1 67.02 75.74 4.46 / 42.56
Heads 2 65.39 72.35 4.47 / 42.92
Heads 3 64.01 71.94 4.47 / 43.28
Varying Number of Attention Heads (Distance = [1,2,3])
Heads 1 60.99 75.38 4.50 / 44.68
Heads 2 69.50 78.12 4.51 / 45.06
Heads 3 65.87 75.87 4.51 / 45.42
Heads 4 61.01 72.98 4.52 / 45.78
Varying Temporal Window Size (Heads = 2)
Distance [1,2] 60.78 71.10 4.48 / 43.98
Distance [1,2,3] 69.50 78.12 4.51 / 45.06
Distance [1–4] 64.18 71.59 4.53 / 46.12
Distance [1–5] 67.54 67.41 4.55 / 47.18
Distance [1–6] 64.03 75.38 4.57 / 48.26
Varying Clip Length with Distance=[1,2,3], Heads=2
Length 25 56.34 64.85 4.47 / 10.72
Length 50 58.56 66.80 4.47 / 21.46
Length 100 69.50 78.12 4.47 / 47.52
Length 150 62.24 74.50 4.47 / 64.22
Length 200 57.29 66.78 4.47 / 85.50
4.5 Ablation Study
We conduct ablation studies on the TTA and FineDiving datasets, which represent two contrasting types of sports—one with high event density and the other with sparse events.
Isolating Components. We evaluate the effectiveness of our proposed extensions to GSM by isolating them individually. First, we test the multi-scale GSM with different temporal distance ranges. As shown in Table 3, increasing the range alone leads to inconsistent performance, since exchanging more temporal information without guidance can introduce noise. Next, we assess the channel-group spatial attention GSM by applying different numbers of attention heads to the baseline E2E-GSM. A single head provides a slight gain by directing focus to the most salient region, but additional heads do not yield consistent improvements. This suggests that attention alone lacks sufficient temporal context to fully exploit multiple regions of interest.
Finally, when combining both components, their individual limitations are mitigated: channel-group spatial attention highlights the most informative regions, while multi-scale shifts propagate this focused information across time. Together, they form MSAGSM, which consistently outperforms the baseline.
Attention Heads and Temporal Window Size. We study the effect of varying attention heads (with distance fixed at $[1,2,3]$) and dilation ranges (with heads fixed at two). On TTA, increasing the number of heads from one to two substantially improves mAP (+8.51 and +2.74 at $\delta=1,2$), but adding more heads leads to inconsistent or reduced performance, likely due to over-fragmentation.
For distance ranges, the default $[1,2,3]$ provides the strongest and most stable results. Slight gains are seen with $[1,2,3,4]$, but very large windows (e.g., $1,\dots,6]$) dilute fine-grained cues and hurt stability. Overall, while optimal hyperparameters may vary slightly with dataset characteristics, two attention heads and a temporal window of $[1,2,3]$ yield the most consistent performance across benchmarks on the validation sets. All results in the main tables are reported using these default settings, with further ablations including validations are provided in the Supp. §3.3.
Clip length. We also examine the impact of clip length. As shown in Table 3, a length of 100 frames achieves the best performance, which is consistent with prior findings in [12].
5 Limitations and Future Work
MSAGSM enables rich temporal modeling with minimal overhead, making it a practical choice for real-world applications. We observe that it performs particularly well with both strong backbones and lightweight models, suggesting that better feature representations further amplify its effectiveness.
Although our work focuses on PES, MSAGSM could also benefit broader video understanding tasks such as action recognition, where sparse frame sampling limits temporal precision. Future work may explore how fine-grained modules like MSAGSM can bridge this gap.
We acknowledge two key limitations. First, MSAGSM is sensitive to hyperparameters (e.g., temporal distance, attention heads), requiring dataset-specific tuning and reducing plug-and-play generalization across datasets. Second, per-class precision–recall analysis on TTA (Supp. §3.6) shows frequent confusion between forehand and backhand strokes due to their visual similarity and player handedness, whereas distinct events like bounces are recognized more reliably.
Future directions include adaptive strategies for automatically learning temporal distances and methods to better handle fine-grained ambiguities and viewpoint variations.
6 Conclusion
We propose MSAGSM, a lightweight temporal module for 2D backbones that combines temporal shifts and attention. Across five fine-grained PES datasets, it consistently improves performance, boosting mAP@0 by +4.09 and mAP@1 by +3.08% on TTA with only 0.01M extra parameters. We also present TTA, a new table tennis PES benchmark with realistic recordings and high event density, to support future research in fine-grained sports video understanding.
Appendix A Notation Table
For clarity, we provide a summary of the main symbols used throughout the paper in Table 4. This table lists the definitions of key variables (scalars, vectors, matrices, and tensors) along with evaluation parameters such as class weights and temporal tolerance, to ensure consistent interpretation of our mathematical notation.
Table 4: Summary of notation used in the paper.
Symbol Description
$t$ Frame index
$L$ Clip length (number of frames per input)
$x_{t}$ Input frame at time $t$
$F_{t}\in\mathbb{R}^{C\times H\times W}$ Feature map at frame $t$ ($C$ channels, $H\times W$ spatial size)
$C,H,W$ Number of channels, height, and width of feature maps
$X_{\text{left}},X_{\text{right}}$ Partitioned input features for left and right shifts
$G_{\text{left}},G_{\text{right}}$ Learned gating maps (via 3D convolution + $\tanh$)
$\odot$ Element-wise (Hadamard) product
$\operatorname{Roll}(X,\pm\delta)$ Circular shift (temporal roll) of features by $\delta$ frames left/right
$\delta$ Temporal shift distance (in frames)
$\operatorname{Concat}(\cdot,\cdot)$ Concatenation of feature groups along the channel dimension
$\hat{X}$ Updated features after gated shifts and concatenation
$\hat{X}^{(i)}$ Output from the $i$-th temporal scale
$w_{i}$ Learnable weight for scale $i$ ($w_{i}\geq 0$, $\sum_{i}w_{i}=1$)
$\hat{X}_{\text{final}}$ Weighted sum of multi-scale shifted features
$y_{t},\hat{y}_{t}$ Ground-truth event label and predicted score at frame $t$
$w$ Class weighting factor in cross-entropy loss
$\delta$ (eval) Temporal tolerance used in evaluation metrics
Appendix B Implementation Details
B.1 Backbone Model
In addition to the backbone model RegNetY-200 [25] used in all three original works—E2E-Spot [12], T-DEED [44], and UGLF [37]—we also experimented with ResNet-18 [10] as the backbone. MSAGSM is integrated into the backbone at each residual block.
B.2 Baselines
For all implementations, we adopted the publicly released code of each baseline whenever available and used the exact training configurations provided in their papers. We did observe minor performance drops compared to the results reported in the original papers, but these were marginal. Since our module is integrated directly into their official implementations without altering other components, we consider this a fair comparison.
We included ASFormer [50] as a transformer baseline, following prior work [12, 44, 27]. In earlier studies [12], ASFormer was evaluated using TSP as the feature extractor, pretrained on video action recognition datasets and then finetuned on PES. To ensure fairness, we instead used RegNetY as the feature extractor for both ASFormer and ASTRA [43], since ASTRA was originally designed to operate on pre-extracted SoccerNet features, including audio inputs.
Regarding UGLF [37], their paper utilizes additional GLIP-provided information, which is not publicly available. Therefore, we followed their released code without this input.
B.3 Training Configuration
All experiments were conducted on the same machine with a single L40S GPU. For fairness, we closely follow the training protocol of E2E-Spot [12]. Each epoch samples 5,000 clips to reduce training time while maintaining coverage of the dataset, and event frames are weighted 5× in the loss to address class imbalance. We train for 50 epochs using AdamW with an initial learning rate of $1\times 10^{-3}$, a batch size of 8, and 3 warm-up epochs followed by cosine decay. As in E2E-Spot, cross-entropy loss is applied, with inputs shaped $[B,T,C,H,W]$ (batch, time, channels, height, width) and outputs $[B,T,C]$, representing frame-wise class probabilities after softmax.
B.4 Data Augmentations
During training, we randomly apply color jitter, Gaussian blur, and mixup [52]. On all datasets, frames of size $398\times 224$ are randomly cropped to $224\times 224$ pixels. The crop is applied only along the width dimension to preserve critical spatial information, as cropping the height could exclude important event regions. This also improves processing speed, since both the tennis court and table tennis table occupy the central region of the frame.
Figure 7: Examples of occluded events, when events happened during occlusion.
Figure 8: Examples from the TTA dataset. Each class is illustrated, with the top row showing examples from Paralympics matches and the bottom row from WPF tournaments.
Table 5: Performance comparison (mAP %) under different temporal tolerances ($\delta=0,1,2$) for T-DEED and UGLF under ResNet-18 backbone. Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp TTA Tennis FS_Comp FS_Perf FineDiving P (M) / GFLOPs
$\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2
E2E-Spot ResNet-18 GSM 40.30 69.13 78.37 49.74 90.43 96.72 37.80 77.31 87.52 43.73 82.75 94.56 26.53 64.63 84.79 14.35 / 367.04
E2E-Spot ResNet-18 GSF 39.11 66.42 80.75 49.32 90.48 97.01 35.50 73.21 87.37 39.04 78.50 90.58 24.31 64.25 84.01 14.35 / 367.04
E2E-Spot ResNet-18 Ours 43.81 70.49 79.38 52.82 90.81 97.52 40.49 78.29 90.93 45.75 85.23 96.62 30.43 67.17 84.95 14.37 / 370.40
T-DEED ResNet-18 GSM 39.81 66.78 76.13 48.09 89.46 97.19 38.64 76.65 87.02 42.82 83.67 95.35 25.86 63.11 83.34 37.34 / 369.26
T-DEED ResNet-18 GSF 38.26 68.84 75.59 45.57 88.48 97.22 35.54 75.58 88.87 42.03 83.09 96.22 28.81 68.68 85.21 37.34 / 369.26
T-DEED ResNet-18 Ours 39.51 69.92 77.33 51.56 87.18 95.85 36.89 76.17 88.57 40.90 82.58 96.35 23.68 66.78 84.88 37.37 / 372.62
UGLF ResNet-18 GSM 32.89 60.05 69.50 49.05 89.38 97.14 36.57 75.93 90.16 42.30 81.28 92.60 24.53 65.88 84.92 14.35 / 367.04
UGLF ResNet-18 GSF 36.98 62.58 68.57 47.59 89.11 97.31 34.22 74.88 88.49 41.73 82.34 94.27 22.42 65.56 84.56 14.35 /367.04
UGLF ResNet-18 Ours 37.14 62.30 72.14 52.14 91.06 97.28 40.69 77.21 90.25 40.55 81.46 94.84 29.56 68.98 87.60 14.37 / 370.40
Appendix C TTA Dataset
We introduce the Table Tennis Australia (TTA) dataset, the first benchmark dataset for table tennis in the PES task.
All videos in the TTA dataset were recorded and provided by the Paralympics Table Tennis Australia team. The dataset is intended strictly for research purposes and will be available upon request and approval.
It consists of 39 games, where each game is an 11-point set and each match is played as best-of-five games. We ensured diversity by including videos captured from different camera angles, with varying table and stadium colors, and across different competition levels—from elite Paralympics events to slightly lower-level WPF tournaments, which reflect broader public scenarios.
The dataset contains annotations for 8 action classes: serve, forehand, backhand, and bounce, each labeled separately for the close table and far table. Examples of each class are shown in Figure 8.
C.1 Dataset Annotation Details
Annotation was performed by three team members, with all annotations subsequently reviewed and verified by a professional sports analyst from Table Tennis Australia. For each event type, we annotated the exact frame corresponding to the critical action—for example, the frame when the ball contacts the bat for a serve, or when it lands on the table (accompanied by a sudden directional change) for a bounce.
We note that due to the frame rate of the videos, some critical moments might not be captured precisely as shown in Figure 7; in such cases, the closest possible frame was annotated. Another challenge arose from the consistent back-facing camera angle: certain events were occasionally blocked or occluded by players—such as close-table bounces or backhands. In these cases, following the procedure used by professional sports analysts, we inferred the event frame based on the surrounding context in adjacent frames.
Appendix D Additional Experiments
In this section, we provide additional experiments using ResNet-18 as the backbone within the E2E, T-DEED, and UGLF frameworks. We also report mAP@0 results across different temporal modules and datasets. Finally, we include results for alternative dilation configurations that were not shown in the main paper.
Table 6: Ablation study on attention head count and dilation range using the MSAGSM module on the TTA and FineDiving datasets in the Validation Set. All results are reported in mAP (%). We also report the average performance across the two datasets.
Setting Value TTA FineDiving Average P (M) / GFLOPs
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2 Avg@0 Avg@1 Avg@2
Varying Number of Attention Heads (Dilations = [1,2,3])
Heads 1 41.97 68.08 72.26 31.72 73.03 88.28 36.85 70.56 80.27 4.50 / 44.68
Heads 2 38.78 71.67 79.37 33.72 73.58 89.77 36.25 72.63 84.57 4.51 / 45.06
Heads 3 38.48 67.23 71.59 30.68 71.29 86.32 34.58 69.26 78.96 4.51 / 45.42
Heads 4 40.13 67.20 70.87 32.98 74.81 87.81 36.56 71.01 79.34 4.52 / 45.78
Heads 8 39.91 66.61 73.54 30.26 72.67 88.20 35.09 69.64 80.87 4.55 / 47.24
Varying Temporal Window Size (Heads = 2)
Distance [1,2] 34.07 64.46 70.68 31.63 74.08 87.63 32.85 69.27 79.16 4.48 / 43.98
Distance [1,2,3] 38.78 71.67 79.37 33.72 73.58 89.77 36.25 72.63 84.57 4.51 / 45.06
Distance [1,…,4] 37.54 65.48 71.84 30.95 72.72 86.47 34.25 69.10 79.16 4.53 / 46.12
Distance [1,…,5] 42.45 69.47 74.16 31.77 74.19 88.41 37.11 71.83 81.29 4.55 / 47.18
Distance [1,…,6] 36.19 67.87 74.91 32.37 74.11 87.79 34.28 70.99 81.35 4.57 / 48.26
Distance [1,…,9] 45.20 74.37 79.10 32.78 70.21 85.37 38.99 72.29 82.24 4.63 / 51.46
D.1 Different Backbone
We evaluate MSAGSM with ResNet-18 as the backbone to test its effectiveness with a different feature extractor. Results in Table 5 show that MSAGSM consistently improves performance across all models, confirming its robustness to backbone choice. Out of 10 results, MSAGSM achieves 8 best scores, and across 30 groups, it achieves 20 best results—particularly excelling in the simpler E2E and UGLF models. As observed in the main results, the additional temporal information from MSAGSM may conflict with the more complex design of T-DEED, leading to less optimal gains in that setting. This suggests that T-DEED’s existing fine-grained temporal modules already capture much of the relevant temporal context, leaving little room for MSAGSM to contribute further. In some cases, the overlapping temporal mechanisms may introduce redundancy or even conflicting signals, which could hinder effective feature learning.
D.2 Exact Frame Prediction
In PES settings, mAP@0 measures whether the prediction lands exactly on the ground-truth frame. However, this is often ambiguous even for human annotators, so mAP@0 is considered a supplementary result. These results are presented in Table 5 for ResNet-18.
We observe that even under the strict mAP@0 setting, MSAGSM achieves strong results: on ResNet-18, in 10 out of 15 cases. This suggests that the module is particularly effective at pinpointing exact event moments when integrated into simpler architectures such as E2E-Spot, as it can fully leverage its multi-scale shifts and spatial attention. In contrast, in more complex models like T-DEED—which already incorporate advanced temporal mechanisms—MSAGSM may partially overlap or conflict with existing components, slightly reducing its relative effectiveness.
Table 7: Sensitivity analysis of the cross-entropy (CE) loss weighting factor on the TTA and FineDiving datasets. All results are reported in mAP (%).
Loss / Weight TTA FineDiving
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2
CE (1) 29.65 58.42 69.56 31.61 65.13 85.75
CE (3) 32.04 60.52 73.38 27.09 67.08 83.16
CE (5) 36.43 69.50 78.12 29.83 71.03 87.26
CE (7) 35.50 67.83 77.15 29.76 68.97 86.63
CE (9) 30.76 60.09 73.40 32.77 69.43 86.61
Focal 32.54 58.65 71.72 30.71 67.93 85.94
Table 8: Ablation study on attention head count and dilation range using the MSAGSM module on the TTA and FineDiving datasets. All results are reported in mAP (%).
Setting Value TTA FineDiving P (M) / GFLOPs
mAP@0 mAP@1 mAP@2 mAP@0 mAP@1 mAP@2
Varying Number of Attention Heads (Dilations = [1,2,3])
Heads 1 32.90 60.99 75.38 29.91 66.36 83.47 4.50 / 44.68
Heads 2 36.43 69.50 78.12 29.83 71.03 87.26 4.51 / 45.06
Heads 3 35.98 65.87 75.87 28.68 67.93 85.11 4.51 / 45.42
Heads 4 30.34 61.01 72.98 30.19 68.72 85.16 4.52 / 45.78
Heads 8 35.26 64.19 72.94 29.03 70.06 85.91 4.55 / 47.24
Varying Temporal Window Size (Heads = 2)
Distance [1,2] 29.64 60.78 71.10 30.18 68.56 85.49 4.48 / 43.98
Distance [1,2,3] 36.43 69.50 78.12 29.83 71.03 87.26 4.51 / 45.06
Distance [1,3,5] 29.83 68.14 75.26 29.69 68.18 85.26 4.51 / 45.06
Distance [1,…,4] 30.06 64.18 71.59 28.85 69.02 85.17 4.53 / 46.12
Distance [1,2,4,6] 39.77 68.27 75.07 32.68 69.85 87.02 4.53 / 46.12
Distance [1,…,5] 29.41 67.54 67.41 30.20 69.26 85.26 4.55 / 47.18
Distance [1,…,6] 35.07 64.03 75.38 29.55 69.84 85.34 4.57 / 48.26
Distance [1,…,9] 41.02 67.32 76.44 27.76 65.47 82.80 4.63 / 51.46
D.3 Validation-Based Hyperparameter Selection
We selected our default hyperparameters by running experiments on validation sets from two datasets. As shown in Table 6, while the optimal configuration may vary across datasets due to differing characteristics, the default setting of $[1,2,3]$ with two attention heads provides the most consistent performance with minimal complexity overhead.
D.4 Loss Function Sensitivity Analysis
We conducted a sensitivity analysis of the Cross-Entropy (CE) loss weighting for the foreground (event) class, and also compared CE with focal loss on two datasets (Table 7). We observe that a weighting factor of 5 yields the best performance on both datasets, outperforming focal loss. This aligns with the standard practice reported in prior work such as E2E-Spot [12], T-DEED [44], and UGLF [37].
D.5 Different Temporal Window Setting
Here, we present a more detailed evaluation of temporal window settings for MSAGSM as shown in Table 8, including wider ranges such as $1,\dots,9]$ and skipped patterns like $[1,2,4,6]$. As longer temporal distances are incorporated, the improvements become unstable and often diminish, suggesting that shifting information from nearby frames is sufficient in most cases. Notably, the configuration $[1,2,3]$ consistently performs best, effectively balancing short-term and moderate long-term dependencies. Smaller windows (like $[1,2]$) fail to capture enough context, while larger windows (like $1,\dots,9]$) include redundant or noisy temporal information, which can confuse the model and harm precision. Thus, moderate dilation allows the model to attend to meaningful temporal cues without overfitting to irrelevant frames.
D.6 Class-wise Spotting Performance
The difficulty of precisely spotting events varies across event classes. In Figure 10, we present the interpolated precision-recall curves for the different classes in the TTA, Tennis, Figure Skating, and FineDiving datasets, as predicted by our default MSAGSM-E2E model.
We show both curves at a tolerance of 1 and a tolerance of 2. In the TTA dataset, the model exhibits difficulties with both-sides backhand events, while performing particularly well on bounce events. A similar pattern is observed in the Tennis dataset. This is reasonable, as bounce frames typically exhibit a distinct characteristic of sudden motion change at the ball’s location, making them easier to detect. In contrast, distinguishing between forehand and backhand is often ambiguous because some players are left-handed and others are right-handed, making the visual cues less consistent for the model.
D.7 Visualization of the Model
We visualize attention maps using Grad-CAM to highlight how each model attends to critical regions during precise event spotting as shown in Figure 9. For table-bounce events under occlusion, baselines such as E2E, T-DEED, and UGLF often fail to focus on the relevant regions, whereas our approach consistently attends to the table area, demonstrating its ability to leverage long-range temporal context and effective attention mechanisms. Similarly, for jump-landing events, our model correctly emphasizes the skater’s feet, rather than irrelevant regions.
Figure 9: Grad-CAM visualizations comparing E2E-MSAGSM with baseline models (E2E, T-DEED, and UGLF) across different event types. For close table bounces under occlusion, only our model consistently attends to the correct table region, while baselines often focus elsewhere. For far table bounces, E2E-MSAGSM maintains attention on the bounce area, whereas others drift toward irrelevant regions. For jump-landing events, our model emphasizes the skater’s feet, in contrast to baselines that spread attention across less informative areas. These visualizations demonstrate the effectiveness of our module in capturing both long-range temporal context and spatially discriminative cues.
(a) TTA Curve
(b) Tennis Curve
(c) FS Comp Curve
(d) FS Perf Curve
(e) FineDiving Curve
Figure 10: Precision-Recall curves for different datasets. The left plot shows results at a tolerance of 1, and the right plot at a tolerance of 2.